OpenStudy (samsan9):

What are the coordinates of the center of the circle shown below? x^2+y^2-12x-6y+9=0

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2+y^2-12x-6y+9=0 }\) can we try to do it together this time ?

OpenStudy (samsan9):

yes i am home now sorry about that

OpenStudy (solomonzelman):

It is okay

OpenStudy (solomonzelman):

First, we will say \(\normalsize\color{red}{ x^2-12x+y^2-6y+9=0 }\) what have I done to the equation ?

OpenStudy (samsan9):

you put the x's next to the x^2 and the y next to the y^2 so that we could perfect the squares?

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

But we can already notice a perfect square in our (red) equation, can't we? (If we can) where is it ?

OpenStudy (samsan9):

i think it is the y^2-6y part

OpenStudy (solomonzelman):

Close, I'll underline it. \(\normalsize\color{black}{ x^2-12x+\underline{ y^2-6y+9}=0 }\)

OpenStudy (solomonzelman):

See ?

OpenStudy (samsan9):

yes i see it now

OpenStudy (solomonzelman):

what do you have to add to both sides, (and were do you put it) to make a `perfect square` where the `x`s are ?

OpenStudy (samsan9):

is it 11 and i think its over the = to the right

OpenStudy (solomonzelman):

This is always the rule about a perfect square. \(\normalsize\color{black}{ x^2±bx=c }\) \(\normalsize\color{black}{ x^2±bx\color{red}{+(b/2)^2}=c \color{red}{+(b/2)^2} }\)

OpenStudy (solomonzelman):

in this case, when looking at `x² - 12 ` what is the `(b/2)²` ?

OpenStudy (samsan9):

it is b=-12 and i got 36

OpenStudy (solomonzelman):

So, you add 36 to both sides of the equation?

OpenStudy (solomonzelman):

Right ?

OpenStudy (solomonzelman):

Can you do that for me please ?

OpenStudy (samsan9):

X^2-12x+9=-36

OpenStudy (sidsiddhartha):

hey @samsan9 do u know diffrentiation ?

OpenStudy (solomonzelman):

SAMSON, Hold on, you had \(\normalsize\color{blue}{ x^2-12x+y^2-6y+9=0 }\), and we said that \(\normalsize\color{blue}{ x^2-12x+\color{red}{y^2-6y+9 }=0 }\) and the red is already a perfect square. So, \(\normalsize\color{blue}{ x^2-12x+\color{red}{y^2-6y+9 }=0 }\) and what do you add to the `x` to make the `x`s a perfect square ?

OpenStudy (solomonzelman):

@sidsiddhartha not need in differentiation.

OpenStudy (samsan9):

is it 27?

OpenStudy (sidsiddhartha):

u can solve problems like this easily with partial diff let \[f(x)=x^2+y^2-12x-6y+9\] now diffrentiate it partially with respect to x and equate it to zero\[\frac{ \delta f }{ \delta x }=2x-12\rightarrow 2x-12=0\rightarrow x=6\] and \[\frac{ \delta f }{ \delta y }=2y-6\rightarrow 2y-6=0\rightarrow y=3\] so center is \[(6,3)\] yeah i know @SolomonZelman it's a messy way to do it but it is quite usefull for objective questions because u can calculate centre of any conics using it very fast

OpenStudy (sidsiddhartha):

@samsan9 forget about my method follow solmol :)

OpenStudy (solomonzelman):

I agree, but considering the fact that not everyone learned differentiation.

OpenStudy (sidsiddhartha):

yeah very true

OpenStudy (samsan9):

i am still in precalc so i don't remember learning that in high school

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2-12x+y^2-6y+9=0 }\) \(\normalsize\color{blue}{ x^2-12x\color{red}{+\rm{what?} }+y^2-6y+9=0\color{red}{+\rm{what?} } }\)

OpenStudy (solomonzelman):

samson, can you do this one ?

OpenStudy (samsan9):

x^2-12x+36+y^2-6y+9=0 and then i move the the 36 to the other side so i get x^2-12x+y^2-6y+9=36

OpenStudy (solomonzelman):

you can't just add 36 to one side.

OpenStudy (solomonzelman):

If you add a number you add it to both sides. You got that it is 36 correctly though, so good job ! \(\normalsize\color{blue}{ x^2-12x+y^2-6y+9=0}\) \(\normalsize\color{blue}{ x^2-12x+36+y^2-6y+9=0+36}\) \(\normalsize\color{blue}{ x^2-12x+36+y^2-6y+9=36}\) And now, to finish it up, re-write the `X`s and the `Y`s as perfect squares.

OpenStudy (samsan9):

i got (x^2-12x+36)(x^2-6y+9) and then i got (-9,-3) and ( (-3,-3)

OpenStudy (solomonzelman):

you wrote 2 points `(-9,-3)` and `(-3,-3)` what are these ?

OpenStudy (samsan9):

i got those using the diamond method

OpenStudy (solomonzelman):

Alright.. but can you tell me the center?

OpenStudy (samsan9):

(6,3)

OpenStudy (solomonzelman):

No you are a little bit off

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2-12x+36+y^2-6y+9=36}\) \(\normalsize\color{blue}{ \underline{x^2-12x+36}+ \underline{y^2-6x+9}=36}\) \(\normalsize\color{blue}{ \underline{(x-6)^2}+ \underline{(y-3)^2}=36}\) No.. I am the bad guy, you're good !

OpenStudy (solomonzelman):

Good work :)

OpenStudy (samsan9):

thank you so much for helping me out this long i really appreciate it :D

OpenStudy (solomonzelman):

You welcome, it is my \(\normalsize\color{blue}{ \rm{ P:} }\)\(\normalsize\color{blue}{ leasure~~ ! }\)